40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15428     Entry time: Wed Jun 24 22:33:44 2020
Author: gautam 
Type: Update 
Category: SUS 
Subject: EQ tripped all suspensions 

This earthquake tripped all suspensions and ITMX got stuck. The watchdogs were restored and the stuck optic was released. The IFO was re-aligned, POX/POY and PRMI on carrier locking all work okay.

ELOG V3.1.3-